OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x \[f(x)=x^3+4\]\[g(x)=\sqrt[3]{x-4}\]

OpenStudy (turingtest):

have you tried anything yet?

OpenStudy (anonymous):

yeah i tried plugging it in but i couldn't understand and ended up confusing myself

OpenStudy (turingtest):

can you try to show what you did? ...or at least the first step?

OpenStudy (anonymous):

sorry my computer is acting up what idid was f(g(x))= Cube root of quantity x minus four + 4

OpenStudy (turingtest):

f(g(x)) will have f on the outside, but you have the cube root, which is in g, on the outside

OpenStudy (turingtest):

\[f(g(x))=[g(x)]^3+4=\left(\sqrt[3]{x^2-4}\right)^3+4\]simplify and do the same for g(f(x))

OpenStudy (anonymous):

but how you Confirm that f and g are inverses by showing that

OpenStudy (anonymous):

how does that equal x

OpenStudy (turingtest):

well, what is\[\sqrt{x^2}\] simplified to?

OpenStudy (turingtest):

or rather, \[(\sqrt x)^2\](which simplifies tothe same thing

OpenStudy (anonymous):

it not x^2 but but just x from the problem above that you solve sorry

OpenStudy (turingtest):

There is an x under the cube root, but I am merely illustrating a point with a related example. if you could just bear with me, you'll see where i'm going with this.I will cahnge symbols to avoid confusion: simply\[(\sqrt n)^2\]what do you get?

OpenStudy (turingtest):

simplify*

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i understand

OpenStudy (turingtest):

ok, you're welcome